OpenStudy (anonymous):

expand the logarithmic expression log(4x^3)/y^2(x-1)^5

OpenStudy (anonymous):

\[\log(4x ^{3}/y^2(x-1)^5 ) \]

OpenStudy (aravindg):

First you can use log (a/b)=log a -log b

OpenStudy (anonymous):

so log4x^3 - log y^2(x-1)^5

OpenStudy (aravindg):

Yep

OpenStudy (aravindg):

Now open log(4x^3) in terms of sum of logs

OpenStudy (anonymous):

im not sure is it 3log4x-2logy(x-1)^5

OpenStudy (aravindg):

Nope log 4x^3=log 4+log x^3=2log 2+3log x Similarly do for 2nd term

OpenStudy (anonymous):

ok so 2logy+5log(x-1) ?

OpenStudy (aravindg):

Right :)

OpenStudy (anonymous):

so I put them over each other or subtract the two........... 2log2+3logx-2logy+5log(x-1) ?

OpenStudy (aravindg):

It should be -5log(x-1) and you are done

OpenStudy (anonymous):

how did it come out to that

OpenStudy (anonymous):

@AravindG

OpenStudy (aravindg):

I just subtracted the 2 terms